OpenStudy (anonymous):

A tunnel links three buildings at 2 Zinc. On rainy days, Dara uses the tunnel to go from building A to building B. On sunny days he uses the path. Dara walks between the buildings 8 times a day. How much more does he walk on a rainy day than on a sunny day?

OpenStudy (anonymous):



OpenStudy (anonymous):

Answers are: 1280ft \[8\sqrt{13600}\] \[8(160-\sqrt{13600}\] \[8(190-\sqrt{13600}\] not enough information

OpenStudy (bibby):

Find AB using the pythagorean theorem. is 8AB greater or less than 8(100+60)?

OpenStudy (anonymous):

ok and after the pythagorean theorem?

OpenStudy (bibby):

Well, first find the length of AB.

OpenStudy (anonymous):

116.62

OpenStudy (anonymous):

Then what?

OpenStudy (bibby):

I got that as well. then we have to measure out the lengths he walks. Our information is as follows: *He uses the path 8 times a day. *On rainy days he uses AB *on sunny days he uses 60+100

OpenStudy (anonymous):

so it would be 8(166.62)(60/100)

OpenStudy (bibby):

I'm not sure where that came from. We have: 8(rainy) - 8(sunny) [or 8(sunny) - 8(rainy) if sunny is the larger one]

OpenStudy (anonymous):

This still does not help me understand or solve the equation

OpenStudy (bibby):

The sunny path is 160 units long, right? That means if he uses the path 8 times, he walks 160*8 units on a sunny day

OpenStudy (bibby):

meanwhile, on a rainy day, he walks 8*116.619... units

OpenStudy (bibby):

oh wait we don't need an actual number

OpenStudy (bibby):

well the difference between the two paths is:\[\huge 160-\sqrt{13600}\] do you see why?